OpenStudy (anonymous):

Find an equation of the line that satisfies the given conditions. Through (2, 8); slope 2/3

OpenStudy (anonymous):

y= mx+b y = 2/3x+b 8 = 2*2/3+b b = 8-4/3 = 20/3 y = 2/3x+20/3 or 3y = 2x+20

OpenStudy (anonymous):

its supposed to be in the form of x+y+___=0

OpenStudy (anonymous):

well, then rewrite it in that form. I can't do EVERYTHING for you! :)

OpenStudy (anonymous):

lol is it 2x-3y+20=0

OpenStudy (anonymous):

yes